# Modeling the amplification of epidemic spread by misinformed populations

### v3

@@ Abstract
Understanding how misinformation affects the spread of disease is crucial for public health, especially given recent research indicating that misinformation can increase vaccine hesitancy and discourage vaccine uptake. However, it is difficult to investigate the interaction between misinformation and epidemic outcomes due to the dearth of data-informed holistic epidemic models. Here, we employ an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data. The model allows us to simulate and estimate various scenarios to understand the impact of misinformation on epidemic spreading. Using this model, we present a worst-case scenario in which a heavily misinformed population would result in an additional 14% of the U.S. population becoming infected over the course of the COVID-19 epidemic, compared to a best-case scenario.
Social factors, such as information sharing, play a crucial role in shaping the dynamics and epidemiology of infectious diseases12, 5. For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures34, 4. A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic40, or by social media and the anti-vaccination movement1, 26, 10, 13.
A great deal of work has explored how to model the influence of human behaviour on the spread of infectious diseases53, 25. Here we focus on risky behaviors affecting disease transmission that are associated with misinformed individuals. Misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance45. Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States42, 39. Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments33. This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity17, 47. Proper management of epidemic crises in the modern age thus requires the understanding of the complex relationship between the spread of (mis)information through online social networks and the spread of disease through physical contact networks (Fig. 1).
@@
While previous agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways50, 36, 41, 9, there is a growing need to integrate real-world data to strengthen the connections between simulation results and real-world outcomes. Social media can provide valuable data sources for this purpose6, 51.
We address this challenge by proposing a data-informed epidemic model that takes both the distribution of misinformed individuals and a physical mobility network into account. Using this data, we augment the Susceptible-Infected-Recovered (SIR) model to account for a subpopulation of “misinformed” individuals. We explore how this group can affect the larger, ordinary population using both a mean-field approximation, which assumes all individuals have an equal chance of interacting, and a multi-level agent-based simulation based on two large, data-informed network: a social network where misinformation spreads and a contact network where the disease can propagate. A contact network of approximately 20 million nodes is constructed by leveraging large-scale Twitter data, county-level voting records, and cellphone mobility data. We incorporate theoretically extreme values of the parameter responsible for the epidemic transmission to evaluate best- and worst-case scenarios about the impact of misinformed individuals on the spread of disease and obtain quantitative bounds on the harm caused by misinformation.
@@ Results
In our model, as in the standard SIR modeling framework2, a parameter $\beta$ describes the average number of infected individuals generated by an infected individual in a time unit. Infected individuals recover with rate $\gamma$. We extend this epidemic model to account for misinformed and ordinary subpopulations. We utilize six compartments to capture the three stages of SIR for each subpopulation. The fractions of individuals in the susceptible, infected, and recovered compartments are denoted by $S_{O}$, $I_{O}$, and $R_{O}$ respectively for the ordinary group; and by $S_{M}$, $I_{M}$, and $R_{M}$ for the misinformed group.
Ordinary individuals are considered to be well-informed about public health guidelines, while misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow recommended behaviors, such as social distancing, mask-wearing, and vaccination, thereby increasing the risk of infection for themselves and others20. We therefore define two distinct parameters $\beta_{O}$ and $\beta_{M}$ for ordinary and misinformed susceptibles, respectively. We model the risky behavior of the misinformed group by setting $\beta_{M}\geq\beta_{O}$. The recovery rate $\gamma$ is identical for both subpopulations. We refer to this as the Susceptible Misinformed Infected Recovered (SMIR) model.
Mean-field simulations
The mean-field approach assumes that an infected individual can transmit disease to any susceptible individual. With this assumption, we derive the model’s mean-field system of equations (see Methods) and explore parameter values for which epidemic spread occurs. We denote the proportions of ordinary and misinformed individuals as $\mu$ and $1-\mu$, respectively. An initial proportion of the population $\epsilon$ is infected, split evenly between the ordinary and misinformed groups. Thus, we initialize the compartments with $I_{O}=\frac{\epsilon}{2}$, $I_{M}=\frac{\epsilon}{2}$, $S_{O}=\mu-\frac{\epsilon}{2}$, $S_{M}=1-\mu-\frac{\epsilon}{2}$, and $R_{O}=R_{M}=0$.
The effects of varying the transmission and recovery parameters are predictable (see Supplementary Information): lower $\beta$ and higher $\gamma$ “flatten the curve” and reduce the negative outcomes of an infection. Based on these explorations we set $\beta_{O}=0.3$ and $\gamma=0.2$ such that the basic reproduction number $R_{0}=\beta/\gamma>1$ to ensure epidemic spread. Next we focus on the effects of varying the misinformed subpopulation’s level of risk and the degree of homophily within the network.
Figure 2: Increasing $\lambda=\beta_{M}/\beta_{O}$ accelerates and amplifies the infection. We use $\beta_{O}=0.3$, $\gamma=0.2$, and $\mu=0.5$. (A) Proportion of the population infected on each day. (B) Extra proportion of the total population that is infected as a function of $\lambda$.
To explore the effects of risky behaviors by misinformed individuals, let us assume two equally-sized subpopulations ($\mu=1/2$) and introduce the scaling factor $\lambda=\beta_{M}/\beta_{O}\geq 1$. Fig. 2 illustrates the increasing negative impact of the misinformed subpopulation on the disease-spreading dynamics as $\lambda$ becomes larger. The social cost associated with the more risky behaviors by the misinformed group is passed on to the whole network. For example, when $\lambda=3.0$, peak infection for the entire population is reached 38 days earlier than in the $\lambda=1$ case (day 22 vs. 60; Fig. 2A), leading to an additional 29.2% of the population becoming infected (Fig. 2B).
Figure 3: Homophily in the contact network worsens the infection among misinformed individuals, especially for lower transmission rates. The combined effects of transmission and homophily parameters, $\beta_{O}$ and $\alpha$, are examined with the mean-field approximation when $\lambda=3$, $\gamma=0.2$, and $\mu=0.5$. We plot the proportions of infected individuals in (A) the ordinary population, (B) the misinformed population, and (C) the overall population. The maximum proportion of the overall population infected for each $\beta_{O}$ is marked with a black dot. When the transmission rate is sufficiently high, homophily benefits the entire population but harms the misinformed group.
Fig. 3 illustrates the effects of homophily ($0.5\leq\alpha\leq 1$) for different levels of ordinary transmission ($0.1\leq\beta_{O}\leq 0.4$). For less infectious disease (low $\beta_{O}$), increasing homophily significantly harms the misinformed group (Fig. 3B): the infection remains confined within this group. There is no discernible effect on the ordinary population as long as the two groups interact; when they do not ($\alpha=1$), we observe a sharp reduction in infections (Fig. 3A). For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig. 3C 37. Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig. 3B and C). As $\beta_{O}$ increases further, while nearly the entire misinformed population becomes infected regardless of $\alpha$ (Fig. 3B), high homophily shields the full population (Fig. 3C): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals. In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 46, 49.
Agent-based simulations
@@
To examine the influence of misinformation within a more realistic epidemic framework, we utilize a multi-level model. Our approach combines an empirically derived information network with a contact network calibrated with real-world data, as illustrated in Fig. 4. Information diffusion is modeled by leveraging a large set of users of a popular social media platform and epidemic simulations are subsequently conducted on contact networks populated with misinformed individuals.
@@
We leverage a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines22. From approximately nine months of this data (Jan. 4–Sep. 30, 2021), we geolocate over 2 million U.S. users who shared almost 26 million tweets and focus on accounts in 341 U.S. counties containing more than 200 Twitter users each. We also infer an account’s political alignment and whether they shared any likely misinformation (see Methods). Twitter is not representative of the U.S. population, and people also access information in other ways, such as traditional media and word of mouth. However, this empirical social network serves as a data-informed proxy for a large, realistic network through which people share information about the disease.
With this data, we build a directed and weighted information diffusion network, in which an edge $(i\rightarrow j,w)$ indicates that $j$ retweeted $i$ $w$ times. There are various ways to model the infodemic19. We simulate the spread of misinformation on this network, as illustrated in Fig. 4A. Accounts that share or reshare posts containing misinformation are considered misinformed. These accounts function as the initial seeds from which misinformation proliferates to the wider network. Not all users actively participate in content sharing, with only about half of U.S. Twitter users, for instance, engaging in this activity38. But even without active sharing, exposure to misinformation can still influence individual behavior33.
To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process27. While many social influence models have been proposed14, this is a simple way to capture complex contagion, according to which individuals require multiple exposures to misinformation before they become misinformed themselves16, 54, 35. Let a linear threshold $\phi$ represent the minimum number of misinformed friends needed for an ordinary node to become misinformed. If the total number of misinformed friends of $i$ is greater than or equal to $\phi$, $i$ is marked as misinformed ($M$). The remaining nodes are marked as ordinary susceptibles ($O$). We can interpret $\phi$ as a measure of “resilience” to misinformation; as it increases, individuals require more exposure to misinformation to be converted to the misinformed group. Note that since we explore the full range of $\phi$ values, the following results are unaffected whether the threshold is defined based on the number of users or the number of retweets.
@@
We generate contact networks for different thresholds ($1\leq\phi\leq 20$) to compare the impact of misinformed subpopulations of different sizes. Given a threshold $\phi$ and the corresponding information network, we aim to construct a physical contact network containing empirically calibrated misinformed subpopulations (Fig. 4B). The process begins by selecting a sample of individuals from each county within the information network. As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic29, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S. presidential election. For each county, we add the sampled nodes to the physical network marked as misinformed ($M$) or ordinary susceptible ($O$), based on their label within the retweet network. Sampling with replacement allows us to select individuals such that the overall proportions of Republicans and Democrats match the voting records. A 10% sample leads to $N\approx 20$ million nodes. A network based on a much smaller sample is illustrated in Fig. 1B. This process captures empirical measurements of the ideological split, relative population size, and quantity of misinformed individuals in each county. We add contact network edges by leveraging cell phone mobility data that provides the probability of an individual traveling within and between counties. See Methods for details.
@@
Disease-spreading dynamics on a contact network are simulated using an agent-based version of SMIR (Fig. 4C). Two critical parameters affecting these dynamics are the density of the contact network, captured by its average degree $\bar{k}$, and the transmission probability $p$ (note that $p\bar{k}=\beta$). Mitigation measures such as social distancing decrease $\bar{k}$, while those such as masking and vaccination decrease $p$. A simple way to model the combined effects of misinformation on these behaviors is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic32, and explore a wide range of transmission probabilities to capture worst- and best-case scenarios. To this end, we model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals. In contrast, we model the adoption of several mitigation measures by selecting an extremely small value $p_{O}=0.01$ for ordinary individuals. These values illustrate two theoretically extreme scenarios. The former portrays a realistic number of interactions during non-pandemic times, accompanied by high transmission rates due to the absence of preventive measures, such as social distancing, mask-wearing, or vaccinations. The latter demonstrates decreased daily interactions and reduced transmission rates resulting from the implementation of these preventive measures. Using the empirically calibrated contact networks in conjunction with these extreme parameters, the simulation approach allows us to bound the best- and worst-case scenarios in a data-informed manner (see Methods for more information).
@@
The worst-case effect of the misinformed subpopulation size on the daily incidence of infection (illustrated in Fig. 1C-F on a small network) is quantified in Fig. 5B on a large network (10% sample). A heavily misinformed population ($\phi=1$) corresponds to an additional 9% of the population being infected at peak time (a six-fold increase) compared to a population following expert guidance in the best-case scenario ($\phi=20$). The peak also occurs approximately two weeks earlier. The cumulative effect of a heavily misinformed population is also significant, with an additional 14% of the population infected over the course of the epidemic — a 32% relative increase (Fig. 5C). We repeated this analysis for different sample sizes and found that the main results are robust (see Supplementary Information).
@@ Discussion
This work presents large-scale simulations based on an epidemic model that accounts for a subpopulation of misinformed individuals, informed by real-world data. Such an approach answers calls from the literature to anchor models in empirical data6, 51, allowing us to explore realistic potential outcomes. While we know that exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal39, our model further links these behaviors by a misinformed population to an accelerated spread of disease.
By deriving the mean-field system of equations for the novel SMIR model and exploring its dynamics, we demonstrate how the risky behaviors of misinformed individuals can adversely impact those following public health guidelines, worsening outcomes for the entire population. Additionally, increasing homophily can benefit the overall population by protecting ordinary citizens; however, it may also lead to higher infection rates within the misinformed subpopulation.
Agent-based simulations of the SMIR model let us study the epidemic on empirically calibrated contact networks, which provide more realistic scenarios compared to the mean-field approach. The model estimates that, in the worst-case scenario, a heavily misinformed population can amplify the peak of the infection by a factor of six and accelerate it by two weeks. This would result in an additional 14% of the population becoming infected — nearly 47 million Americans based on recent U.S. Census data52. The corresponding price tag of vaccine misinformation would be over \$143B, using estimated health care costs associated with COVID-19 in the U.S.3.
@@
We acknowledge certain limitations in our approach. While individual beliefs and behaviors may vary over time, our model simplifies the scenario by dichotomizing individuals into misinformed and ordinary subpopulations and assuming constant transmission rates. We also assume uniform resilience to misinformation for all individuals during the information diffusion process, although this attribute likely differs across individuals. Future directions could involve more sophisticated models to account for these heterogeneities. For instance, cognitive models of misinformation acceptance7 could be incorporated into the simulation with misinformation exposure data collected from social media. Such integration would enable the transition of individuals from ordinary to misinformed susceptible states throughout the simulation, allowing for a simultaneous examination of opinion and disease dynamics. Some theoretical models have already explored similar approaches and obtained results that align with our findings36, 50.
Future work should attempt to analyze the potential effects of misinformation in more realistic scenarios, where values for the key parameters $p_{M}$ and $p_{O}$ could be obtained by calibrating the model with empirical epidemic data from particular regions and time periods. Additional extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population.

### v4

# Modeling the amplification of epidemic spread by individuals exposed to misinformation on social media

@@ Abstract
Understanding how misinformation affects the spread of disease is crucial for public health, especially given recent research indicating that misinformation can increase vaccine hesitancy and discourage vaccine uptake. However, it is difficult to investigate the interaction between misinformation and epidemic outcomes due to the dearth of data-informed holistic epidemic models. Here, we employ an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data. The model allows us to simulate various scenarios to understand how epidemic spreading can be affected by misinformation spreading through one particular social media platform. Using this model, we compare a worst-case scenario, in which individuals become misinformed after a single exposure to low-credibility content, to a best-case scenario where the population is highly resilient to misinformation. We estimate the additional portion of the U.S. population that would become infected over the course of the COVID-19 epidemic in the worst-case scenario. This work can provide policymakers with insights about the potential harms of exposure to online vaccine misinformation.
Social factors, such as information sharing, play a crucial role in shaping the dynamics and epidemiology of infectious diseases12, 5. For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures36, 4. A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic42, or by social media and the anti-vaccination movement1, 27, 10, 14.
A great deal of work has explored how to model the influence of human behavior on the spread of infectious diseases59, 26. Here we focus on risky behaviors affecting disease transmission that are associated with misinformed individuals. Misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance48. Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States45, 41. Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments35. This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity18, 50. Proper management of epidemic crises in the modern age thus requires the understanding of the complex relationship between the spread of (mis)information through online social networks and the spread of disease through physical contact networks (Fig. 1).
@@
Agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways54, 38, 44, 9. One model estimated that between March and November 2021, misinformation caused at least 198 thousand additional COVID-19 cases, 2,800 additional deaths, and \$299M in additional hospital costs in Canada30. However, there is a growing need to strengthen the connections between simulation results and real-world outcomes by integrating real-world data from social media6, 55.
We address this challenge by proposing a data-informed epidemic model that takes both the distribution of misinformed individuals and a physical mobility network into account. Using this data, we augment the Susceptible Infected Recovered (SIR) model to account for a subpopulation of “misinformed” individuals. We refer to this as the Susceptible Misinformed Infected Recovered (SMIR) model. We explore how the misinformed group can affect the larger, ordinary population using a multi-level agent-based simulation based on two large, data-informed networks: a social network where misinformation spreads and a contact network where the disease can propagate. A contact network of approximately 20 million nodes is constructed by leveraging large-scale Twitter data, county-level voting records, and cell phone mobility data. We incorporate theoretically extreme values of the parameter responsible for the epidemic transmission to evaluate best- and worst-case scenarios about the impact of misinformed individuals on the spread of disease and obtain quantitative bounds on the harm caused by misinformation. The proposed model lets us move beyond simplified experimental settings to assess the impacts of misinformation56.
@@ Results
We utilize a multi-level, agent-based model to examine the influence of misinformation on epidemic spread. Our approach combines an empirically derived information network with a contact network calibrated with real-world data, as illustrated in Fig. 2. Information diffusion is modeled by leveraging a large set of users of a popular social media platform. Epidemic simulations are subsequently conducted on contact networks populated with misinformed individuals.
@@
We start from a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines23. From approximately nine months of this data (Jan. 4–Sep. 30, 2021), we geolocate over 2 million U.S. users who shared almost 26 million tweets and focus on accounts in 341 U.S. counties containing more than 200 Twitter users each. We also infer an account’s political alignment and whether they shared any likely misinformation (see Methods). Twitter is not representative of the U.S. population, and people also access information in other ways, such as traditional media and word of mouth. However, this social media platform serves as one large, realistic network through which people share information about the disease.
With this data, we build a directed and weighted information diffusion network, in which an edge $(i\rightarrow j,w)$ indicates that $j$ retweeted $i$ $w$ times. There are various ways to model the infodemic20. We simulate the spread of misinformation on this network, as illustrated in Fig. 2A. Accounts that share or reshare posts containing misinformation are considered misinformed. These accounts serve as the initial seeds from which misinformation proliferates, with exposure to this content likely concentrated within the wider network13. Many users may not actively participate in content sharing; for instance, only about half of U.S. Twitter users engage in sharing40. Even without active sharing, exposure to misinformation or misleading content can still influence individual behavior35, 1.
To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process28. While many social influence models have been proposed15, this is a simple way to capture complex contagion, according to which individuals may require multiple exposures to misinformation before they become misinformed themselves17, 60, 37. Let a linear threshold $\phi$ represent the minimum number of misinformed friends needed for an ordinary node to become misinformed. If the total number of misinformed friends of $i$ is greater than or equal to $\phi$, $i$ is marked as misinformed ($M$). The remaining nodes are marked as ordinary susceptibles ($O$). We can interpret $\phi$ as a measure of “resilience” to misinformation; as it increases, individuals require more exposure to misinformation to be converted to the misinformed group. Conversely, we can think of $\phi$ as inversely related to intent or motivation to engage with low-credibility content52. Note that since we explore the full range of $\phi$ values, the following results are unaffected whether the threshold is defined based on the number of users or the number of retweets.
@@
We generate contact networks for different thresholds ($1\leq\phi\leq 20$) to compare the impact of misinformed subpopulations of different sizes. Given a threshold $\phi$ and the corresponding information network, we aim to construct a physical contact network containing empirically calibrated misinformed subpopulations (Fig. 2B). The process begins by selecting a sample of individuals from each county within the information network. As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic31, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S. presidential election. For each county, we add the sampled nodes to the physical network marked as misinformed ($M$) or ordinary susceptible ($O$), based on their label within the retweet network. Sampling with replacement allows us to select individuals such that the overall proportions of Republicans and Democrats match the voting records. A 10% sample leads to $N\approx 20$ million nodes. A network based on a much smaller sample is illustrated in Fig. 1B. This process captures empirical measurements of the ideological split, relative population size, and quantity of misinformed individuals in each county. It also allows us to account for the known link between the ideological motivations of users and their exposure to misinformation13, 41. We add contact network edges by leveraging cell phone mobility data that provides the probability of an individual traveling within and between counties. See Methods for details.
Disease-spreading dynamics on the contact network are simulated using the SMIR model (Fig. 2C). As in the standard SIR2, a parameter $\beta$ describes the average number of infected individuals generated by an infected individual in a time unit. We can express $\beta=p\bar{k}$ in terms of two critical parameters that affect the spreading dynamics: the density of the contact network, captured by its average degree $\bar{k}$, and the transmission probability $p$. Infected individuals recover with rate $\gamma$.
We extend this epidemic model to account for misinformed and ordinary subpopulations. Ordinary individuals are considered to be well-informed about public health guidelines, such as social distancing, mask-wearing, and vaccination. Mitigation measures such as social distancing decrease $\bar{k}$, while those such as masking and vaccination decrease $p$. Misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow these recommended behaviors, thereby increasing the risk of infection for themselves and others21. A simple way to model the combined effects of misinformation on these behaviors through a single parameter is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic34, and use extreme values of $p$ to capture worst- and best-case scenarios. An effective reduction of contacts, resulting for example from social distancing or lockdowns, can be captured by decreasing the $p$ parameter.
@@
We therefore model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals. In contrast, we model the adoption of several mitigation measures by selecting an extremely small value $p_{O}=0.01$ for ordinary individuals. The former scenario portrays a realistic number of interactions during non-pandemic times, accompanied by high transmission rates due to the absence of preventive measures, such as social distancing, mask-wearing, or vaccinations. The latter demonstrates decreased daily interactions and reduced transmission rates resulting from the implementation of these preventive measures. Using the empirically calibrated contact networks in conjunction with these extreme parameters, the simulation approach allows us to bound the best- and worst-case scenarios in a data-informed manner (see Methods for more information).
@@
The effects of the misinformed subpopulation size on the daily incidence of infection (illustrated in Fig. 1C-F on a small network) are quantified in Fig. 3B on a large network (10% sample). The worst case capturing a heavily misinformed population ($\phi=1$) corresponds to an additional 9% of the population being infected at peak time (a six-fold increase) compared to a resilient population following expert guidance in the best-case scenario ($\phi=20$). The peak also occurs approximately two weeks earlier. The cumulative effect is also significant, with an additional 14% of the population infected over the course of the epidemic compared to case with a more resilient population — a 32% relative increase (Fig. 3C).
We explored alternative scenarios for the ratio $p_{M}/p_{O}$ through a mean-field approximation. Predictably, as this ratio gets larger, the infected population increases and the peak infection occurs earlier. We also considered different sample sizes for the empirical network and found that the main results are robust. These analyses can be found in Supplementary Information.
@@ Discussion
Exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal41. There is also experimental evidence suggesting a causal link35, 57, 1. While one study found no evidence that misinformation reduces intent to vaccinate, the authors report that they did not have sufficient power to detect small effects43. Assuming an association exists between exposure to health misinformation on one particular social media platform and risky behaviors, this work uses large-scale epidemic simulations to further link the behaviors of misinformed individuals to an accelerated spread of disease. Our model is anchored in empirical data6, 55 to explore potential outcomes.
Agent-based simulations of the SMIR model let us study the epidemic on empirically calibrated contact networks. By comparing a worst-case scenario, in which individuals become misinformed after a single exposure to low-credibility content, to a best-case scenario where the population is highly resilient to misinformation, the model estimates that the peak of the infection is amplified by a factor of six and accelerated by two weeks. This would result in an additional 14% of the population becoming infected — nearly 47 million Americans based on recent U.S. Census data58. The corresponding price tag of vaccine misinformation would be over \$143B, using estimated health care costs associated with COVID-19 in the U.S.3.
@@
Our results do not address the differential effects of the epidemic on the two populations of ordinary and misinformed individuals. We carry out such an analysis using a mean-field approximation of the model, which assumes all individuals have an equal chance of interacting. The mean-field model demonstrates how the risky behaviors of misinformed individuals can adversely impact those following public health guidelines, worsening outcomes for the entire population (see Supplementary Information). Additionally, we use the mean-field model to explore the role of homophily in the population, i.e., scenarios where misinformed individuals are more likely to be connected to other misinformed individuals and similarly for the ordinary population. We find that increasing homophily can benefit the overall population by protecting ordinary citizens; however, it may also lead to higher infection rates within the misinformed subpopulation (see Supplementary Information).
We acknowledge several limitations in our approach. The model assumes the existence of a causal link between exposure to online misinformation and the adoption of risky behaviors. There is a need for models that can provide support for this assumption beyond existing lab experiments35, 1.
Using empirical retweet data as a proxy for social connections may not capture potential passive exposure to misinformation. While follower relationships could diminish this limitation, our choice allows us to focus on users who are more likely to be impacted due to their active engagement.
We model a single wave of infection with somewhat arbitrary extreme-case parameters ($p_{O}=0.01,p_{M}=1$). A broader range of values is explored in a mean-field scenario, along with the effect of the size of the misinformed population (see Supplementary Information). Of course, as $p_{O}\rightarrow 0$, only the misinformed population can get infected in the model. However, since the mean-field scenario ignores the network structure, its results cannot be directly compared to those of the agent-based model. COVID-19 saw multiple waves of infection with different variants, varying reproduction numbers, levels of immunity, and so on. Future work should attempt to quantify the potential effects of misinformation in more realistic scenarios, where the key parameters $p_{M}$ and $p_{O}$ could be calibrated on empirical surveillance data from particular regions and time periods.
We also assume uniform resilience to misinformation for all individuals during the information diffusion process, although this attribute likely differs across individuals. Future directions could involve more sophisticated models to account for these heterogeneities. For instance, cognitive models of misinformation acceptance7 could be incorporated into the simulation with misinformation exposure data collected from social media. Such integration would enable the transition of individuals from ordinary to misinformed susceptible states throughout the simulation, allowing for a simultaneous examination of opinion and disease dynamics. Some theoretical models have already explored similar approaches and obtained results that align with our findings38, 54.
Finally, although individual beliefs and behaviors may vary over time, our model simplifies the scenario by dichotomizing individuals into misinformed and ordinary subpopulations and assuming constant transmission rates. Future extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population62.